Please help me to do this sum.
The population of the town increases every year by 2 % of the population at the beginning of that year.The number of years by which the total increase of population be 40% is
Please as I am getting wrong answer ,So please help me
ThankYou



Answer by chessace(471)

Let x be the number of years it takes to reach +40%.
Let P(0) = initial population, P(1) = After 1st year, etc.
So we need to find x such that P(x) = 1.4 P(0).
P(1) = 1.02 * P(0)
P(2) = 1.02 * P(1) = 1.02 * (1.02 * p(0)) = (1.02)^2 P(0)
...
P(x) = (1.02)^x P(0)
(1.02)^x = 1.4.
x * log(1.02) = Log(1.4)
x = log(1.4) / log(1.02)
x = 0.3365 / 0.0198 = 17 years